Find the mass and the center of mass of a triangular lamina with vertices (0.0), (1,0) and (0,2) if the density function is p=1+3x+y


1
Expert's answer
2020-10-27T19:17:45-0400

Mass,

"M=\int_0^1\int_0^2(1+3x+y)\,dy\,dx\\
=\int_0^1[y+3xy+\frac{y^2}{2}]_0^2dx\\
=\int_0^1(6x+4)dx=[3x^2+4x]_0^1=7"

Thus, mass is 7 unit.

Now,


"X_{com}=\frac{1}{M}\int_0^1\int_0^2x(1+3x+y)dydx\\
=\frac{1}{M}\int_0^1(6x^2+4x)dx\\
=4/7"

"Y_{com}=\frac{1}{M}\int_0^1\int_0^2y(1+3x+y)dydx\\
=\frac{1}{M}\int_0^1(6x+\frac{14}{3})dx\\
=23/21"